# MuLan: Adapting Multilingual Diffusion Models for Hundreds of Languages with Negligible Cost

### v1

@@ 4.3 Qualitative Results
Our model can generate high-quality images across multiple languages. It also supports multilingual mixed input and can recognize certain emojis. In addition to simple T2I tasks, there are numerous downstream applications of Stable Diffusion (SD) within the community, including LoRA [13], LCM [23], and ControlNet [49]. These tools played crucial roles in enhancing model adaptability, control over outputs, and finetuning for specific tasks. Unlike AltDiffusion [46], our model does not require finetuning on SD, allowing it to seamlessly integrate and be compatible with these community-developed SD applications in a plug-and-play manner. We show some examples in Figure 4.
@@ 5 Conclusion
We introduce language adapter MuLan that could equip image/video/3D diffusion models with multilingual generation abilities. MuLan shows strong zero-shot capabilities for up to 110 different languages, even if the adapter is solely trained on English data. MuLan also can be trained with a frozen text encoder and diffusion denoising model, which makes it applicable for many downstream models, such as LoRA [13], ControlNet [49], LCM [23], and etc., without any additional finetuning. MuLan is currently trained with paired data, and it could inevitably bring in bias and cause a distribution shift of original models. A promising extension would be to alleviate the need for paired data and make original capabilities intact. Furthermore, MuLan currently focuses on improving multilingual generation capabilities, but it would be interesting to extend it to improving prompt understanding and following under a multilingual context.
[…]
Supplementary Material
@@ Appendix A More Experiments
In this section, we provide supplementary analyses to further evaluate and understand the performance and efficiency of our proposed model. These studies delve into key aspects such as training efficiency, adapter architecture, semantic alignment methods, and the model’s generalization across multiple languages. By examining the impact of various design choices and alignment strategies, we aim to shed light on the model’s multilingual capabilities and resource efficiency.
@@ A.1 Training Efficiency
$\lambda$ of samples Mean CLIP Sim
@@
Table 8: Impact of dataset size on multilingual performance of the model. We evaluated the model’s performance using different proportions of the original dataset, ranging from full size ($\lambda=1$) to $1/1024$, and measured the Mean CLIP similarity scores across 7 languages. The results show that the model maintains relatively strong performance even with significantly reduced data sizes, demonstrating its efficiency and robustness under resource-constrained settings.
@@
Setting. We choose Stable Diffusion 1.5 [33] as the backbone model, using the AdamW optimizer for training. The model was trained to convergence with a learning rate of 2e-5 and a batch size of 128, without employing any training tricks. Additionally, the dataset size was reduced, and we set multiple data proportion levels ranging from 0.5 to 0.001 to evaluate the impact of data size on the model’s performance.
@@
Results. As shown in Table 8, we found that the model’s multilingual performance decreased as the data volume was reduced; however, it still maintained a relatively strong performance until we reduced the data to $1/1024$ of the default data volume. This level of data volume and training cost is highly developer-friendly, requiring only 48 GPU hours to achieve decent multilingual text-to-image (T2I) capabilities for the model. Compared to existing multilingual T2I models, such as AltDiffusion [46], our approach requires significantly less data and computational cost at every stage.
A.2 PixArt-$\alpha$ with MuLan
We also trained MuLan combined with Pixart-$\alpha$, which we called InternVL-MuLan-Pixart-$\alpha$. We evaluated InternVL-MuLan-Pixart-$\alpha$ on the Crossmodal3600 dataset, and the model demonstrated strong multilingual text-to-image (T2I) capabilities, performing exceptionally well in common languages. This also confirms that our approach is effective not only for Diffusion models with a UNet [34] architecture but also for models based on the DiT [27] architecture.
en ar ja zh fr
Ours 39.52 39.12 40.38 39.28 40.17
AD18 39.98 38.33 39.37 38.98 40.01
Table 9: CLIP similarity scores of InternVL-MuLan-Pixart-$\alpha$ and AD18 (AltDiffusion-m18 [46]) on crossmodal3600 [41]. Our model outperforms across multiple languages.
A.3 Comparison of Semantic Alignment Methods
Here, we conducted a comparison of two semantic alignment methods to explore the impact of different semantic alignment approaches on multilingual image generation.
Setting. We selected XLM-RoBERTa-Large[7] as the text encoder for comparison.
(1) For language-centered alignment, we conducted distillation training following Multilang-CLIP [3]. We utilized a larger translation dataset, CCMatrix [37], consisting of 2.91B samples. The choice of a large translation dataset was made to match the scale of LAION-5B [36], enabling a comparison between the effectiveness of translation data and image-text data in aligning multilingual representations. We selected CLIP text encoder as the teacher model and XLM-RoBERTa-Large [7] as the student model. The training process is configured with a batch size of 512 and a constant learning rate of 2e-5, applied over 10 epochs.
(2) For image-centered alignment, we selected ViT-XLM-R-L [16] provided by LAION and InternVL-LLaMA [6] as the text encoders. These text encoders were trained on the LAION-5B [36] dataset and aligned with ViT using contrastive learning.
Model Align Method en es fr zh
LLaMA2-7B [42]
XLM-R-L [7]
CCMatrix XLM-R-L [3] Language-centered 36.3 34.2 35.1 33.2
LAION5B XLM-R-L [16] Image-centered 36.5 36.7 37.2 34.5
InternVL-LLaMA [6] Image-centered 38.6 37.7 38.0 37.7
Table 10: Text Encoder Clip Similarity Scores Across Different Languages. CCMatrix XLM-R-L refers to XLM-RoBERTa-large pre-trained using CCMatrix [37] dataset with distillation learning. LAION5B XLM-R-L refers to CLIP-ViT-H-14-frozen-xlm-roberta-large-laion5B-s13B-b90k [16].
Results. The results are shown in Table 10. We can observe that although translation data is easier to collect and more efficient for training compared to text-image pairs, the performance achieved through text-centered alignment using translation data is inferior to that of image-centered alignment using text-image pairs. This highlights the critical importance of text-image pairs in image generation tasks.
Considering XLM-RoBERTa-Large [7] and other unaligned text encoders, it is evident that under our default experimental setup, training a single language adapter alone is ineffective. Even when trained on English text-image pairs, the model performs poorly when using English prompts and fails to generalize to other languages. However, after the model is aligned through either language-centered or image-centered approaches, it can gain multilingual text-to-image generation capabilities via MuLan.
Id SD15 SD21 SDXL AD18 Id SD15 SD21 SDXL AD18 Id SD15 SD21 SDXL AD18
@@ A.3 Comparison of Semantic Alignment Methods
Table 11: CLIP Sim on the COCO2014 validation set. ‘Id’ indicates the key of different language. SD15/21/XL [33, 28] represent our model implemented on the corresponding three T2I backbones, while AD18 refers to the AltDiffusion-m18 [46]. Our model also demonstrates strong text-to-image (T2I) capabilities across a wider range of languages.
A.4 Detailed Results on COCO2014 validation set
In Section 4.2, we evaluated InternVL-Mulan-SD15 on the COCO2014 [20] validation set (85 languages) and compared it with AltDiffusion-m18 [46]. More results are shown in Table 11. It can be observed that our model generalizes well to a wider range of languages and delivers impressive performance.
A.5 Impact of Semantic Alignment on Multilingual Features
In this section, we conduct a preliminary analysis of the multilingual output features produced by various text encoders to reveal the effects of the two image alignment training methods introduced in Section 3.2. The text encoders we selected for this analysis include XLM-RoBERTa-Large [7], CCMatrix pre-trained XLM-RoBERTa-Large in A.3, ViT-XLM-R-L* (LAION-5B [36] pre-trained XLM-RoBERTa-Large [7]), LLaMA2-7B [42], and InternVL-LLaMA [6].
We employ t-SNE, a nonlinear dimensionality reduction technique, to visualize the aggregation effects of multilingual features produced by these text encoders. t-SNE is particularly suited for preserving the local structure of high-dimensional data in a low-dimensional space. Ideally, for text encoders trained with multilingual alignment methods, the features corresponding to semantically equivalent prompts in different languages should exhibit aggregation effects when projected into a lower-dimensional space.
For this analysis, we randomly sampled 20 captions from the COCO2014 [20] validation set and translated them into 8 languages using machine translation, resulting in a total of 160 textual inputs. These inputs were then encoded using the selected text encoders to obtain their corresponding pooled feature representations.
Figure 5: t-SNE analysis on embeddings of 9 prompts in 20 languages produced by 5 text encoders. (a) XLM-RoBERTa-Large [7] (b) CCMatrix pre-trained XLM-RoBERTa-Large [7] in A.3 (c) Laion-5B [36] pre-trained XLM-RoBERTa-Large [7] (d) LLaMA2-7B [42] (e) InternVL-LLaMA [6]. Points of the same color represent embeddings corresponding to the same prompt translated into different languages.
The results are shown in Figure 5. As shown in Figures (a) and (d), the sample points in the t-SNE [43] plots exhibit a scattered distribution for text encoders that have not undergone image alignment training. This indicates that although XLM-RoBERTa [7] and LLaMA2 [42] possess multilingual representation capabilities, their output features for synonymous multilingual prompts are not closely located in the Euclidean space. In contrast, as shown in Figures (b)(c)(e), the sample points exhibit a clustered distribution after alignment training with images, with each cluster corresponding to the 20 different language versions of a single prompt. Through alignment training with images, the model minimizes the semantic discrepancies between multilingual representations, ensuring that the embeddings of synonymous prompts converge in the vector space.
@@ B.2 Plug and Play on Different Visual Generator
Figure 7: IP-Adapter Results. Our model enables multilingual style transfer by integrating with the IP-Adapter [47].
Our model also supports some existing tools based on the Stable Diffusion series developed by the community. Here, we showcase several popular models, including LoRA [13] models, ControlNet [49] models, and IP-Adapter [47] models. The model’s multilingual capabilities naturally come into play in these applications.
Figure 8: Lora Results. Our model can naturally support multilingual input when using LoRA [13].
Figure 9: ControlNet Results. Our model can utilize existing ControlNet [49] models, enabling multilingual image generation with conditional inputs such as depth maps and pose images.

### v2

@@ 4.3 Qualitative Results
Our model can generate high-quality images across multiple languages. It also supports multilingual mixed input and can recognize certain emojis. In addition to simple T2I tasks, there are numerous downstream applications of SD within the community, including LoRA (Hu et al., 2021), LCM (Luo et al., 2023), and ControlNet (Zhang et al., 2023). These tools played crucial roles in enhancing model adaptability, control over outputs, and finetuning for specific tasks. Our model does not require finetuning on SD, allowing it to seamlessly integrate and be compatible with these community-developed SD applications in a plug-and-play manner. We show some examples in Figure 4.
@@ 5 Conclusion
We introduce language adapter MuLan that could equip image/video/3D diffusion models with multilingual generation abilities. MuLan shows strong zero-shot capabilities for up to 110 different languages, even if the adapter is solely trained on English data. MuLan also can be trained with a frozen text encoder and diffusion denoising model, which makes it applicable for many downstream models, such as LoRA (Hu et al., 2021), ControlNet (Zhang et al., 2023), LCM (Luo et al., 2023), and etc., without any additional finetuning. MuLan is currently trained with paired data, and it could inevitably bring in bias and cause a distribution shift of original models. A promising extension would be to alleviate the need for paired data and make original capabilities intact. Furthermore, MuLan currently focuses on improving multilingual generation capabilities, but it would be interesting to extend it to improving prompt understanding and following under a multilingual context.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A More Experiments
In this section, we provide supplementary analyses to further evaluate and understand the performance and efficiency of our proposed model. These studies delve into key aspects such as training efficiency, semantic alignment methods, and the model’s generalization across multiple languages. By examining the impact of various design choices and alignment strategies, we aim to shed light on the model’s multilingual capabilities and resource efficiency.
@@ A.1 Training Efficiency
$\lambda$ of samples Avg SIM
@@
Table 7: Impact of dataset size on multilingual performance of the model. We evaluated the model’s performance using different proportions of the original dataset, ranging from full size ($\lambda=1$) to $1/1024$, and measured the average CLIP Score across 7 languages. The results show that the model maintains relatively strong performance even with significantly reduced data sizes, demonstrating its efficiency and robustness under resource-constrained settings.
@@
Setting. We choose Stable Diffusion 1.5 (Rombach et al., 2022) as the backbone model, using the AdamW optimizer for training. The model was trained to convergence with a learning rate of 2e-5 and a batch size of 128, without employing any training tricks. Additionally, the dataset size was reduced, and we set multiple data proportion levels ranging from 0.5 to 0.001 to evaluate the impact of data size on the model’s performance.
@@
Results. As shown in Table 7, we found that the model’s multilingual performance decreased as the data volume was reduced; however, it still maintained a relatively strong performance until we reduced the data to $1/1024$ of the default data volume. This level of data volume and training cost is highly developer-friendly, requiring only 48 GPU hours to achieve decent multilingual text-to-image (T2I) capabilities for the model. Compared to existing multilingual T2I models, such as AltDiffusion (Ye et al., 2023a), our approach requires significantly less data and computational cost at every stage.
A.2 Impact of Semantic Alignment on Multilingual Features
In this section, we conduct a preliminary analysis of the multilingual output features produced by various text encoders to reveal the effects of the two image alignment training methods introduced in Section 3.2. The text encoders we selected for this analysis include XLM-R Large (Conneau, 2019), CCMatrix trained XLM-R Large* in Table 6, Mul-OpenCLIP (LAION-5B (Schuhmann et al., 2022) pre-trained XLM-RoBERTa-Large (Conneau, 2019)), LLaMA2-7B (Touvron et al., 2023), and InternVL-LLaMA (Chen et al., 2023b).
We employ t-SNE (van der Maaten & Hinton, 2008), a nonlinear dimensionality reduction technique, to visualize the aggregation effects of multilingual features produced by these text encoders. t-SNE is particularly suited for preserving the local structure of high-dimensional data in a low-dimensional space. Ideally, for text encoders trained with multilingual alignment methods, the features corresponding to semantically equivalent prompts in different languages should exhibit aggregation effects when projected into a lower-dimensional space.
For this analysis, we randomly sampled 20 captions from the COCO2014 (Lin et al., 2015) validation set and translated them into 8 languages using machine translation, resulting in a total of 160 textual inputs. These inputs were then encoded using the selected text encoders to obtain their corresponding pooled feature representations.
Figure 5: t-SNE analysis on embeddings of 9 prompts in 20 languages produced by 5 text encoders. (a) XLM-RoBERTa-Large (Conneau, 2019) (b) CCMatrix pre-trained XLM-RoBERTa-Large (Conneau, 2019) in Table 6 (c) Laion-5B (Schuhmann et al., 2022) pre-trained XLM-RoBERTa-Large (Conneau, 2019) (d) LLaMA2-7B (Touvron et al., 2023) (e) InternVL-LLaMA (Chen et al., 2023b). Points of the same color represent embeddings corresponding to the same prompt translated into different languages.
The results are shown in Figure 5. As shown in Figures (a) and (d), the sample points in the t-SNE (van der Maaten & Hinton, 2008) plots exhibit a scattered distribution for text encoders that have not undergone image alignment training. This indicates that although XLM-RoBERTa (Conneau, 2019) and LLaMA2 (Touvron et al., 2023) possess multilingual representation capabilities, their output features for synonymous multilingual prompts are not closely located in the Euclidean space. In contrast, as shown in Figures (b)(c)(e), the sample points exhibit a clustered distribution after alignment training with images, with each cluster corresponding to the 20 different language versions of a single prompt. Through alignment training with images, the model minimizes the semantic discrepancies between multilingual representations, ensuring that the embeddings of synonymous prompts converge in the vector space.
Id SD15 SD21 SDXL AD Id SD15 SD21 SDXL AD Id SD15 SD21 SDXL AD
@@ A.2 Impact of Semantic Alignment on Multilingual Features
Table 8: CLIP Score on the COCO2014 validation set. ‘Id’ indicates the key of different language. SD15/21/XL (Rombach et al., 2022; Podell et al., 2023) represent our model implemented on the corresponding three T2I backbones, while AD refers to the AltDiffusion (Ye et al., 2023a). Our model also demonstrates strong text-to-image (T2I) capabilities across a wider range of languages.
A.3 Detailed Results on COCO2014 validation set
In Section 4.2, we evaluated InternVL-Mulan-SD15 on the COCO2014 (Lin et al., 2015) validation set (85 languages) and compared it with AltDiffusion (Ye et al., 2023a). More results are shown in Table 8. It can be observed that our model generalizes well to a wider range of languages and delivers impressive performance.
@@ B.2 Plug and Play on Different Visual Generator
Figure 7: IP-Adapter Results. Our model enables multilingual style transfer by integrating with the IP-Adapter (Ye et al., 2023b).
Our model also supports some existing tools based on the Stable Diffusion series developed by the community. Here, we showcase several popular models, including LoRA (Hu et al., 2021) models, ControlNet (Zhang et al., 2023) models, and IP-Adapter (Ye et al., 2023b) models. The model’s multilingual capabilities naturally come into play in these applications.
Figure 8: Lora Results. Our model can naturally support multilingual input when using LoRA (Hu et al., 2021).
Figure 9: ControlNet Results. Our model can utilize existing ControlNet (Zhang et al., 2023) models, enabling multilingual image generation with conditional inputs such as depth maps and pose images.